Use the following functions that show quantity demanded and quantity supplied to answer the question. Qd = 100 – 5P and Qs = –20 + 3P. If the price is R10, then there will be a shortage of

1
Expert's answer
2021-06-03T18:40:04-0400

At equilibrium, quantity demanded = quantity supplied

100 - 5P = -20 + 3P

8P = 120

Pe = 15

If equilibrium price = R15; equilibrium quantity Q = 100 - 5(15)

Qe = 25

At price R10: Quantity demanded is 100 - 5(10) = 50

Quantity supplied is -20 + 3(10) = 10

If the price = R10, there will be a shortage of:

10 - 50 = -40

There'll be a shortage of 40